5. A household buys 30 units of a commodity when its price is ₹5 per unit. The quantity demanded 

falls to 25 units when the price rises to ₹6 per unit. How much is price elasticity of demand. 


1
Expert's answer
2021-09-13T11:22:10-0400

Price elasticity of demand is given by

"=\frac{\frac{Q_2-Q_1}{\frac{Q_2+Q_1}{2}}}{\frac{P_2-P_1}{\frac{P_2+P_1}{2}}}"


"=\frac{\frac{25-30}{\frac{25+30}{2}}}{\frac{6-5}{\frac{6+5}{2}}}"


"=\frac{\frac{5}{27.5}}{\frac{1}{55}}\\=\frac{-0.182}{0.182}\\=-1"